OpenStudy (anonymous):

I need help with 3 geometry problems please!!

OpenStudy (sweetburger):

what are the questions?

OpenStudy (anonymous):



hartnn (hartnn):

so where are you stuck?! \(6^2 =36 \\ 8^2 =64\) add them!

OpenStudy (anonymous):

add 36 and 64 together?

hartnn (hartnn):

yes that will give you \(x^2 \)

hartnn (hartnn):

so what will be 'x' ?

OpenStudy (anonymous):

100

hartnn (hartnn):

that will be x^2 \(x^2 =100 \\ x = \sqrt{100} = .. ?\)

OpenStudy (anonymous):

10

hartnn (hartnn):

correct! 10 just like in your username ;)

OpenStudy (anonymous):

Lol! :)

hartnn (hartnn):

Can you ask other questions in another post? :)

OpenStudy (anonymous):

ok!